OpenStudy (anonymous):

find the exact area of the region [-2,10] from f(x)=2x-4 with 6 sub intervals using Riemann Sums and sigma notation.

OpenStudy (anonymous):

What type of reimann sum? left, right or middle?

OpenStudy (anonymous):

the exact one. \[\sum_{i=1}^{n}\] c =cn. it is the old way on integrating

OpenStudy (anonymous):

c=cn, i=(n(n+1))/2, i^2 = (n(n+1)(2n+1))/6